OpenStudy (anonymous):

Use the angle difference identities to find the exact value of the trigonometric expression. cos 195° can someone tell me the answer pleasee?

OpenStudy (anonymous):

this is what the question is asking you to do, to use the cos of the diffrence of angles...

OpenStudy (anonymous):

i'm sorry i dont understand?

OpenStudy (ash2326):

Use A=240 B=45

OpenStudy (anonymous):

okay thank you :) one second.

OpenStudy (ash2326):

Use this \[\cos (A-B)=\cos A \times \cos B+\sin A \times \sin B\]

OpenStudy (anonymous):

how do i get sin A and sin B?

OpenStudy (anonymous):

just sub the values that @ash produces

OpenStudy (anonymous):

this is the formula for cos(A-B)

OpenStudy (anonymous):

do you want a proof for that formula...

OpenStudy (ash2326):

\[\sin 240=\sin (180+60)=-\sin 60\]

OpenStudy (anonymous):

i get 195 when i do Cos(A-B)

OpenStudy (ash2326):

Yes, but you need to find the value of \(\cos 195\) So we use \[\cos (A-B)\] and then we expand it to find the value of cos 195

OpenStudy (anonymous):

240-45

OpenStudy (anonymous):

im not following. i did (240 -45) and got 195

OpenStudy (anonymous):

i don't understand what i'm supposed to do , this is supposed to be an Algebra II course but its trig problems i dont even know what cos and sin is.

OpenStudy (ash2326):

You haven't studied trigonometry, have you?

OpenStudy (anonymous):

not at all.

OpenStudy (ash2326):

Oh, then how would you know this? Was this part of your course?

OpenStudy (anonymous):

yes , this is my last course before i graduate and it says nothing about any of this in my textbooks so im at a loss.

OpenStudy (anonymous):